OpenStudy (anonymous):

what is the equivalent fraction of 84/112

OpenStudy (anonymous):

what you can do is simplify 84=2*2*3*7 112=2*2*2*2*7 find the common multiples so 84=4*7*3 112=4*7*4 \[\frac{ 84 }{ 112 }\div \frac{ 28 }{ 28 }=\frac{ 3 }{ 4 }\]